OpenStudy (anonymous):

If logr 6 = m and logr 3 = n, then logr is equal to

OpenStudy (anonymous):

If \[\log _{r} 6 = m, \log _{r} 3 = n\], then \[\log _{r} (r/2)\] is equal to?

jimthompson5910 (jim_thompson5910):

First, notice that 6/3 = 2. So this means that if we subtract the given equations, we get \[\large m-n=\log_{r}(6)-\log{r}(3)=\log_{r}\left(\frac{6}{3}\right)=\log_{r}(2)\] So \[\large m-n=\log_{r}(2)\] ---------------------------------------------------- Now let's simplify the given expression \[\large \log_{r}\left(\frac{r}{2}\right)\] \[\large \log_{r}\left(r\right)-\log_{r}\left(2\right)\] \[\large 1-\log_{r}\left(2\right)\] \[\large 1-(m-n)\] \[\large 1-m+n\] ------------------------------------------- This means that \[\large \log_{r}\left(\frac{r}{2}\right)=1-m+n\]

OpenStudy (anonymous):

So for any logarithmic expression in which the base is the same as its factor the expression can be reduced to 1?

OpenStudy (anonymous):

And thanks, by the way. The answer was correct

jimthompson5910 (jim_thompson5910):

Exactly, for any number b (where b is not 1 and is nonnegative), the following is true \[\large \log_{b}(b)=1\]

OpenStudy (anonymous):

Ah, I don't know that. Thanks!